1. How many MOLECULES of phosphorus pentachloride are present in 3.40 moles of this compound?   

2. How many MOLES of phosphorus pentachloride are present in 7.67×1022 molecules of this compound?
Answer #1

1)
use:
number of molecules = number of mol * Avogadro’s number
number of molecules = 3.4 * 6.022*10^23 molecules
number of molecules = 2.047*10^24 molecules
Answer: 2.05*10^24 molecules

2)
use:
number of mol = number of molecules / Avogadro’s number
number of mol = 7.67*10^22 / 6.022*10^23 mol
number of mol = 0.1274 mol
Answer: 0.127 mol